The 2018 Canadian internet use survey showed, overall, 94% of Canadians had home internet access, and about 87% of the internet users bought goods or services online, spending $57.4 billion. Suppose 81 internet users with home internet access were randomly and independently sampled. Find the probability that at least 20% of those sampled currently did not buy goods or services online. Round your answers to the nearest ten-thousandth (4 decimals).
Answer #1

P(Bought goods from internet) = 0.94*0.87 = 0.8178

So P(Didn't buy) = 1 - 0.8178 = 0.1822

Since sample size is large, we can appy central limit theorem. So,

P(At least 20%)

= P(p > 0.20)

= P(z > 0.42)

= 0.3391